OpenStudy (anonymous):

Find the volume of a solid. x=2 sqroot of (y) x=0 y=9 about the y axis

OpenStudy (anonymous):

its a 3-D question right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

don't u think there should be a z co=ordinate as well

OpenStudy (dumbcow):

Use the disc method. \[V = \pi \int\limits_{a}^{b} f(y)^{2} dy\] \[f(y) = 2\sqrt{y}\] \[V = \pi \int\limits\limits_{0}^{9} (2\sqrt{y})^{2} dy = \pi \int\limits_{0}^{9} 4ydy\]

OpenStudy (dumbcow):

did that help?

OpenStudy (anonymous):

ohhhh its easy

OpenStudy (anonymous):

yess.. thanks

OpenStudy (dumbcow):

if it was about the x-axis, just use f(x) and integrate w/ respect to x

OpenStudy (anonymous):

yea,, thats waht confuses me... but i am gettting it